# Speech emotion recognition from voice messages recorded in the wild

### v1

@@ Abstract
Emotion datasets used for Speech Emotion Recognition (SER) often contain acted or elicited speech, limiting their applicability in real-world scenarios. In this work, we used the Emotional Voice Messages (EMOVOME) database, including spontaneous voice messages from conversations of 100 Spanish speakers on a messaging app, labeled in continuous and discrete emotions by expert and non-expert annotators. We created speaker-independent SER models using the eGeMAPS features, transformer-based models and their combination. We compared the results with reference databases and analyzed the influence of annotators and gender fairness. The pre-trained Unispeech-L model and its combination with eGeMAPS achieved the highest results, with 61.64% and 55.57% Unweighted Accuracy (UA) for 3-class valence and arousal prediction respectively, a 10% improvement over baseline models. For the emotion categories, 42.58% UA was obtained. EMOVOME performed lower than the acted RAVDESS database. The elicited IEMOCAP database also outperformed EMOVOME in the prediction of emotion categories, while similar results were obtained in valence and arousal. Additionally, EMOVOME outcomes varied with annotator labels, showing superior results and better fairness when combining expert and non-expert annotations. This study significantly contributes to the evaluation of SER models in real-life situations, advancing in the development of applications for analyzing spontaneous voice messages.
@@
Speech emotion recognition, natural database, valence, arousal, pre-trained model, speaker independent.
@@ I Introduction
Human communication allows individuals to express not only their ideas, but also their emotional state. In everyday conversations, people utilize both information to adjust their behavior, underscoring the importance of recognizing emotions. Speech Emotion Recognition (SER) is a research field that automatically identifies a person’s emotional state from their voice. SER has potential applications for the study of human-to-human communications, such as detecting stress or depression in medical contexts. It is also relevant in human-computer interactions, for example, improving the naturalness of speech synthesis systems.
A key aspect of SER research is emotional databases containing labeled samples from which to learn patterns correlating with emotions. In the literature, emotions are modeled using continuous or discrete emotion models [1, 2, 3]. The discrete model is based on categories of basic emotions considered innate and universal, which can be combined to obtain other emotions. For example, Ekman [4] proposed the “big six” basic emotions: fear, surprise, happiness, sadness, anger and disgust. The dimensional model explains emotions using continuous dimensions that represent fundamental properties shared by all emotions. Various dimensional models exist, depending on the number and type of dimensions considered. For instance, Russell’s circumplex model of affect [5] uses two dimensions —valence and arousal— on a Cartesian axis system. Valence represents the pleasantness (positive valence) or unpleasantness (negative valence) of emotion, while arousal measures its intensity, ranging from active/exciting (high arousal) to passive/dull (low arousal). Each emotion can be viewed as a combination of these two dimensions, illustrated in Fig. 1. Both emotional models have strengths and weaknesses, as described in [2]. SER literature predominantly focuses on the discrete model due to its intuitive labeling since individuals commonly use basic categories to express emotions in everyday life. However, it struggles to accurately capture some complex affective states. In contrast, the dimensional model is less intuitive, but it provides greater flexibility in categorizing a broader spectrum of emotions. It acknowledges the complexity of emotional experiences, where individuals may experience a mix of emotions rather than fitting into discrete categories, and captures the dynamic by showing how emotions relate to each other on a continuum.
Figure 1: Russell’s circumplex model of affect. The outer circle illustrates where prototypical emotions are normally located. From [6].
Different databases exist depending on whether the samples are acted, elicited/induced and natural/spontaneous speech as described in several reviews [1, 2, 3]. Acted databases comprise speech samples of actors simulating emotions, typically while delivering predetermined sentences. Despite their prevalence in literature, acted emotions tend to be stereotypical and exaggerated, limiting their applicability in real-world scenarios. Elicited databases involve speech samples collected in artificially induced situations to elicit specific emotional states, for instance, listening to a story, watching a video or conducting a guided discussion. While these samples offer a closer approximation to genuine emotional expression, the induction process has limitations and ethical implications. Individuals may react differently to the same stimulus, requiring an extra subjective assessment to identify the emotion in the speech sample. Natural databases include speech samples extracted from diverse sources, including films, radio/TV talk-shows and occasionally therapy interviews. These databases encounter several challenges. They typically contain audios with overlapping voices and background noise, known as in-the-wild conditions. Furthermore, individuals who are aware of being recorded might unintentionally control their emotions or express them unnaturally. Additionally, determining the emotion in each sample necessitates an external subjective evaluation. Natural databases, though infrequent in literature and often private for ethical and legal reasons, are crucial for recognizing genuine, real-life emotions.
To create speech emotion recognition models, different approaches have been studied in the literature. Traditionally, SER systems have leaned on so-called hand-crafted features, i.e., acoustic properties such as prosodic or spectral features extracted at frame level or aggregated using high-level statistics functions [3]. With the advancements in deep learning, it became possible to directly use the raw signal to train the recognition models, capturing the temporal dynamics and avoiding feature engineering. Techniques such as convolutional neural networks and recurrent neural networks emerged as the primary approach in SER [2]. Nevertheless, they have been limited by the small size of the emotional datasets, which is a major challenge for deep learning methods to achieve optimal effectiveness [7]. In recent years, large pre-trained models have emerged as a powerful framework that is gaining significant attention in all speech-related domains, including automatic speech recognition (ASR), speaker verification and SER, among others [8, 3]. These transformer-based models (e.g., Wav2vec 2.0 [9] and HuBERT [10]) are trained using self-supervised learning approaches on large unlabeled datasets through pretext tasks. Then, they are used to extract speech representations, even from long sequences at once, overcoming the limitations of previous deep learning methods. The learned speech representations are fed into simple neural networks in downstream tasks, including SER, in which recent works have obtained promising results [11, 7, 12, 8].
SER is still an open-ended problem due to its complexity. Emotions are subjective internal states, which makes their theoretical conceptualization difficult [13]. Consequently, annotating audios is a challenging task, as emotions can be perceived differently between raters [14], as well as several emotions can also be combined in the same utterance. This is also reflected in the available emotional speech databases in the literature, often constrained by small sizes and limitations in recognizing real-life emotions. The availability of databases is even more limited for languages other than English, such as Spanish. A total of twelve Spanish databases are mentioned in the literature, consisting of six private [15, 16, 17, 18, 19, 20] and six open for research, among which two are acted [21, 22], one elicited [23] and three natural [24, 25, 26]. Furthermore, the performance of the SER task is significantly influenced by the data separation strategy used. Speaker-dependent (SD) models are trained and tested using speech samples from the same speakers, reaching accuracies of over 70% regardless of the databases or emotion classes involved [3]. In contrast, speaker-independent (SI) models use different subjects for training and testing, so they are evaluated with unknown speakers. The majority of SI models yield accuracies ranging from 29% to 65%, indicating the inherent difficulty of SER in these cases [3]. Finally, another challenge for SER models is fairness, that is, the difference in model results with respect to individuals or population groups according to variables such as gender, age or ethnicity [27]. Given the increasing concerns in society about biases in ML systems [28], it is necessary to ensure fairness in SER models.
@@
To fill the gap of existing literature on models trained in real-world settings, this study focuses on speech emotion recognition using a natural speech database we meticulously collected, called the Emotional Voice Messages (EMOVOME) database [29]. To the best of our knowledge, this is the first Spanish database of spontaneous emotions in real voice messages, and it is labeled by expert and non-expert annotators. We created speaker-independent SER models using acoustic features and state-of-the-art pre-trained models. We conducted a comprehensive analysis to assess how various EMOVOME properties influence the results of the SER models, comparing the results with the widely used Interactive Emotional Dyadic Motion Capture (IEMOCAP) database [30] and Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS) [31]. We explored the following Research Questions (RQ):
  • RQ1: What is the performance of SER models in EMOVOME when utilizing classical acoustic features versus state-of-the-art speech embeddings?
  • RQ2: To what extent does the performance of EMOVOME, a natural database, compare with elicited and acted reference databases in the literature, such as IEMOCAP and RAVDESS?
  • RQ3: What impact do annotator labels exert on the performance of SER models when dealing with challenging natural databases like EMOVOME?
  • RQ4: How does gender impact the fairness of SER model outcomes in EMOVOME compared to reference databases?
The rest of the paper is organized as follows. Section 2 provides a review of related works on SER. Section 3 describes the databases used in this study. In Section 4, the methodology applied to create the SER models is detailed, and Section 5 reports the outcomes of these models. Section 6 discusses the results of the stated hypothesis. The paper concludes by summarizing the work and proposing future directions for SER with realistic data.
@@ II Related work
There is a scarcity of research focused on SER using Spanish databases. Table I summarizes the related literature. For each of the twelve Spanish databases found, it indicates the type of speech (acted, elicited or natural), the number of samples (N), the number of speakers (Spk) with the gender distribution (males/females), the labeled emotions and the database access type (private, commercially available or free). It also includes the Unweighted Accuracy (UA) and/or the Weighted Accuracy (WA) of the monolingual SER models created using these databases (results of multilingual and cross-lingual models are not reported here since they represent more challenging tasks and monolingual outperforms the others for comparable number of training samples [32]). Four databases contain mixed languages: the Emotional Speech Synthesis database [22] is part of the INTERFACE project [33], in which data in English, Slovenian, French and Spanish was collected; RekEmozio [20] includes data in both Spanish and Basque; CMU-MOSEAS [25] includes YouTube videos in Spanish, Portuguese, German and French; and EmoFilm [26] includes clips from both original English films and their Spanish and Italian dubbed versions. LIRIS-ACCEDE [34] is also a multilingual database, but it is mostly in English, with only a small subset available in nine other languages (including Spanish). Due to a lack of specific details on these partitions, they have been excluded from the table. Additionally, EmoWisconsin [35] and IESC-child [36] are emotional databases in Spanish; however, they contain children’s speech. As our focus here is on adult speech, they are not included in the table.
TABLE I: Speech emotion recognition models in Spanish. The asterisk (*) indicates if the database is multilingual, but the information corresponds to the Spanish partition. The “big six” basic emotions are: fear, surprise, happiness, sadness, anger, and disgust.
Database Type N Spk (M/F) Emotions Access UA (%) WA (%)
Iriondo et al. [15] Acted 336 8 (4/4) Big six + desire Private - -
Spanish Emotional
Speech [16]
Acted 1288 1 (1/0)
Happy, sad, cold
anger, surprise
Private - -
Martínez & Cruz [17]
Acted +
natural
300 +
80
15 (-/-) +
N/A (-/-)
Happy, sad, anger,
neutral, fear
Private -
55 (tested on natural
partition) [17]
Emotional Mexican
Spanish speech [18]
Acted 240 6 (3/3)
Happy, sad, anger,
neutral
Private $\geq 95$ [18] -
Spanish Expressive
Voices [19]
Acted 3890 2 (1/1)
Big six + neutral
(cold/hot anger)
Private - 95 [19]
RekEmozio * [20] Acted 2618 10 (5/5) Big six + neutral Private - SI: 74.82 [37]
Mexican Emotional
Speech Database [21]
Acted 3456
8 (4/4) +
8 (3/5)
children
Happy, sad, anger,
neutral, fear,
disgust
Free
89.49 female [21]
93.90 male [21]
83.30 children [21]
Emotional speech
synthesis [22] (from
INTERFACE * [33])
Acted 5520 2 (1/1)
Happy, sad, anger,
neutral, fear,
boredom, disgust
Comm.
avail.
90.9 female [38]
89.4 male [38]
@@
 MOUD [24]
Natural
550 105 (21/84)
Positive, neutral,
negative
Free -
46.75 (audio, positive
@@
CMU-MOSEAS * [25]
Natural
10000 341 (-/-)
Big six, sentiment,
subjectivity, others
Free - -
EmoFilm * [26]
Natural
342 57 (33/24)
Happy, sad, angry,
fear, contempt
Free SI: 69.15 [32]
SI: 70.76 [32]
85.0 [42]
91.8 female [42]
97.7 male [42]
Ours (EMOVOME) [29]
Natural
(voice
messages)
999 100 (50/50)
Valence, arousal
Big six + neutral
Free
SI: 49.27% (audio,
3-class valence) [29]
SI: 44.71% (audio,
3-class arousal) [29]
Most of the Spanish emotional databases contain speech acted by professional or non-professional actors [15, 16, 17, 18, 19, 21, 22, 20], one is based on elicited responses [23] and three are natural databases [24, 25, 26]. Six of the databases are private [15, 16, 17, 18, 19, 20], one is commercially available [22] and five are free [23, 24, 21, 25, 26]. This paper focuses on the natural databases and, although the three existing ones are public, they have some limitations. MOUD database [24] comprises 105 video clips manually extracted from YouTube content in Spanish, covering various topics, and with an average duration of 30s each. Two annotators labeled these clips into three categories: positive, neutral, or negative. Notably, the neutral class is represented by only 4 samples, in contrast to 47 samples for the positive class and 54 for the negative class. Additionally, a large proportion of the speakers in the database are female. As for CMU-MOSEAS [25], this database comprises 10000 Spanish sentences extracted from monologue videos on YouTube, with an average sentence duration of 6.7s. Each sentence was annotated by 3 annotators using different labels, including sentiment and the six Ekman’s emotions. The article asserts attempts to achieve gender balance in videos across languages and regions but lacks specific details on the final distribution. Furthermore, despite the public availability of the database, the original clips are not disclosed, and only high-level features are made accessible. Both the MOUD and CMU-MOSEAS databases share two common constraints. Firstly, their speakers record videos for YouTube, and awareness of being recorded might lead to unintentional emotional control or unnatural expression in response to the artificial situation [43]. Secondly, the nature of YouTube videos in both databases involves monologues, potentially lacking the naturalness found in conversational contexts, as highlighted in [44]. Finally, EmoFilm database [26] includes clips from original English films and their Spanish and Italian dubbed versions. The definition of natural databases includes speech extracted from films. However, films inherently involve acting, and as such, the emotions depicted are portrayed by actors. This holds true for the original films and their dubbed versions, potentially rendering the emotional expressions even more overacted in the latter. Therefore, this database may not be suitable for recognizing real emotions, as noted by their authors [26].
The performance of SER models in Spanish databases is conditioned by the type of data used. For acted databases, the accuracy is equal or greater than 90% (e.g. [18, 19, 21, 38, 39, 40, 41]). An exception is the work in [37], in which they used the RekEmozio database to classify seven emotion categories and achieved 74.82% accuracy. However, they used a speaker-independent approach, which is known to yield comparatively lower results [3]. This trend is also reflected in [23], where they used the elicited database EmoSpanishDB to classify seven emotion categories and obtained 56.3% and 42.1% UA for speaker-dependent and speaker-independent, respectively. The difference between both approaches decreased when they used EmoMatchSpanishDB, a refined version of the database obtained by eliminating labeled samples that did not match the original elicited emotion, achieving an accuracy of 65.0% and 64.2% UA for SD and SI, respectively. Finally, natural databases have the lowest scores. Using only the audio modality, in [24] they achieved an accuracy of 46.75% in differentiating between positive and negative samples. In the study by [17], an accuracy of 55% was achieved for the classification of five emotions using a model trained on acted and tested on natural data. Higher scores were achieved with the EmoFilm database, which may be due to the inherent acted nature of the samples. In [32], they obtained 69.15% and 70.76% UA and WA, respectively, using a speaker-independent approach. In [42], the score increased to 85.0% WA, and 91.8% and 97.7% for female and male versions of the models.
@@
Regarding the techniques used to create the SER models, a noteworthy paradigm shift involves leveraging large pre-trained models in emotion recognition tasks. The study in [32] used the pre-trained wav2vec2-large-robust model fine-tuned on MSP-Podcasts dataset from [11] to extract speech embeddings. These embeddings were then fed into a support vector machine to create the models, obtaining 69.15% unweighted accuracy for five emotion categories on the EmoFilm database. To our knowledge, this is the only previous study using pre-trained models for monolingual SER models in Spanish. Nevertheless, an increasing number of recent publications have focused on the use of pre-trained models with English databases (predominantly relying on acted datasets) for categories of emotions and less frequently for dimensions.
@@
As for studies on emotion category prediction, a very recent study [11] summarized the state-of-the-art results for 4-class emotion classification in the widely used IEMOCAP database, including studies using the pre-trained wav2vec 2.0 and HuBERT models. Comparing the cross-validation results, the unweighted accuracy values range from 60.0% to 74.3%, while the weighted average recall values vary from 62.6% to 79.6%, both top models. Overall, HuBERT surpassed the performance of the wav2vec 2.0 models. Other works have explored several pre-trained models and emotional databases. In [45], they used eight pre-trained models (including wav2vec 2.0) and four databases (CREMA-D, TESS, SAVEE, Emo-DB) to predict 6-7 emotions. The results showed that pre-trained models for speaker recognition (x-vector and ECAPA) achieved the highest scores, possibly due to their learned ability to identify unique features within an individual’s speech. Subsequently, UniSpeech-SAT achieved the best results, a model pre-trained using multitask learning, including the speaker identity. In another study [46], the authors compared nineteen pre-trained models (including wav2vec 2.0, HuBERT, UniSpeech-SAT and wavLM) and five datasets (IEMOCAP, MSP-IMPROV, MSP-PODCAST, CMU-MOSEI, JTES) to predict 4-6 emotion categories using a speaker-independent approach. They found that the best scores were achieved with WavLM, UniSpeech-SAT, and HuBERT, all three in the large version.
@@
As for studies on emotion dimension prediction, [7] used a multimodal model (audio + text) as a teacher to fine-tune HuBERT embeddings to predict valence, arousal and dominance on MSP-Podcast database [47]. They obtained state-of-the-art Concordance Correlation Coefficient (CCC) values of 0.757, 0.627 and 0.671 for arousal, valence and dominance, respectively. The previous state-of-the-art performance for valence was 0.377 [48], so 0.627 represents a substantial improvement. They also replicated the results for the IEMOCAP database, achieving again state-of-the-art CCC results for valence (0.667), arousal (0.582) and dominance (0.545). In [11], the authors conducted an exhaustive analysis using different variants of the pre-trained wav2vec 2.0 and HuBERT models for valence, arousal and dominance prediction on the MSP-Podcast database. They obtained the best result in the literature for valence prediction using only audio, with a wav2vec2-large-robust model that achieved a CCC of 0.638. Notably, their results showed that data used for pre-training the models and the fine-tuning of the transformer layers had a strong influence on valence prediction. Both factors played a role in shaping the models’ ability to implicitly incorporate linguistic information embedded in the audio signal. This, in turn, accounts for their success in valence prediction, achieving similar performance to multimodal models that integrate explicit textual information.

### v3

# EMOVOME: A Dataset for Emotion Recognition in Spontaneous Real-Life Speech

@@ Abstract
Spontaneous datasets for Speech Emotion Recognition (SER) are scarce and frequently derived from laboratory environments or staged scenarios, such as TV shows, limiting their application in real-world contexts. We developed and publicly released the Emotional Voice Messages (EMOVOME) dataset, including 999 voice messages from real conversations of 100 Spanish speakers on a messaging app, labeled in continuous and discrete emotions by expert and non-expert annotators. We evaluated speaker-independent SER models using acoustic features as baseline and transformer-based models. We compared the results with reference datasets including acted and elicited speech, and analyzed the influence of annotators and gender fairness. The pre-trained UniSpeech-SAT-Large model achieved the highest results, 61.64% and 55.57% Unweighted Accuracy (UA) for 3-class valence and arousal prediction respectively on EMOVOME, a 10% improvement over baseline models. For the emotion categories, 42.58% UA was obtained. EMOVOME performed lower than the acted RAVDESS dataset. The elicited IEMOCAP dataset also outperformed EMOVOME in predicting emotion categories, while similar results were obtained in valence and arousal. EMOVOME outcomes varied with annotator labels, showing better results and fairness when combining expert and non-expert annotations. This study highlights the gap between controlled and real-life scenarios, supporting further advancements in recognizing genuine emotions.
@@
Speech emotion recognition, spontaneous, valence, arousal, pre-trained model, speaker-independent.
@@ I Introduction
Human communication conveys not only ideas but also emotional states, which play a crucial role in shaping interpersonal interactions. In everyday conversations, individuals rely on both verbal content and emotional cues to interpret meaning and adjust their responses, highlighting the significance of emotion recognition. Speech Emotion Recognition (SER) is an evolving research field aimed at automatically identifying a speaker’s emotional state from their voice. With applications ranging from detecting emotional distress in customer service interactions to enhancing naturalness in speech synthesis systems, SER holds significant promise for improving both human-to-human and human-computer interactions. Central to SER are emotional datasets, which are categorized in the literature based on various criteria, including the emotion models used for labeling, the authenticity of emotional expressions, and the recording environment.
The labeling of emotions in emotional datasets is influenced by the underlying emotion model used, which can be either discrete or continuous. Grounded in basic emotion theories, discrete models consider emotions as universal and distinct categories, including happiness, sadness, and anger among others [1]. In contrast, dimensional models, based on core affect theory, represent emotions as points within a continuous space defined by affective dimensions such as valence and arousal [2]. The choice of an emotion model directly impacts the scope and granularity of the dataset, shaping its suitability for different SER applications.
Emotional datasets are also distinguished by the expression authenticity of their samples. In the literature, they are commonly classified as acted, elicited, and spontaneous [3, 4, 5]. Acted (or simulated) datasets comprise speech samples of actors simulating emotions, typically using predetermined texts. Acted emotions tend to be stereotypical and exaggerated, limiting their applicability in real-life scenarios. Elicited (or induced) datasets gather speech samples from artificial situations designed to elicit specific emotional states, such as listening to a story or watching a video. The induction process has limitations and ethical implications, as individuals may react differently to the same stimulus. Spontaneous (or natural) datasets aim to capture genuine emotional expression that occurs without external guidance or deliberate performance, making them the most challenging to collect.
The recording environment refers to the setting in which emotional data is collected, encompassing the level of control and external influences during the recording process. It can be categorized into staged, laboratory, and real-life settings. Staged settings involve controlled scenarios with carefully designed setups, such as studios for TV shows or YouTube videos, which pose certain limitations. The artificial situation may inadvertently influence speakers, leading to controlled or unnatural emotional expressions [6]. Additionally, the monologue format of some YouTube videos and podcasts may lack the naturalness typically found in conversational interactions [7]. It is also worth noting that the individuals present in these contexts are generally communication experts, limiting the extrapolation to the general population. Laboratory settings provide controlled environments with some degree of spontaneity when used to simulate real-life scenarios, although they can also influence the authenticity of expressions [6]. Exceptions to these limitations exist in datasets from real-life scenarios, such as call center records. These datasets prioritize ecological validity but are typically restricted and come with challenges such as background noise and variability in recording conditions, called in-the-wild conditions.
SER is still an open-ended problem due to its complexity, particularly when addressing spontaneous emotions in real-life environments. Collecting spontaneous emotional data is inherently difficult and relies on subjective external evaluation to identify the emotion in each sample, a process complicated by the variability in emotional perception among raters [8]. Furthermore, real-life settings are scarce in literature and often restricted from public use due to ethical and legal considerations. As a result, while most emotional corpora are predominantly in English, owing to its global prevalence [4], the availability of natural datasets remains limited. Three exceptions including spontaneous emotions from real-life scenarios were found in the literature: NATURAL [9], Lee & Narayanan [10] and SUSAS [11] datasets. The first two include call-center recordings but have restricted public access. SUSAS includes a commercially available subset, but it is limited to recordings of isolated words from aircraft communications. The availability of emotional datasets is considerably limited for languages other than English, including Spanish. We found only two spontaneous datasets in Spanish: MOUD [12] and CMU-MOSEAS [13]. Although both datasets are publicly available, they consist of monologue clips sourced from YouTube, thus recorded in staged controlled settings. The limited availability of natural datasets from real-life scenarios restricts the advancement of SER models in real-world settings.
@@
To fill the gap in existing literature on emotional datasets with spontaneous emotions from real-life scenarios, we collected the Emotional Voice Messages (EMOVOME) dataset. It contains 999 audio messages collected from real WhatsApp conversations of 100 Spanish speakers. To the best of our knowledge, this is the first public dataset with spontaneous emotions from conversations collected in a real-life scenario. We created SER models using acoustic features and speech embeddings derived from state-of-the-art pre-trained models. To simulate more realistic and challenging conditions, we assessed the models under a speaker-independent approach, using different speakers for training and evaluation. We conducted a comprehensive analysis to assess how various EMOVOME properties impact on SER model results. We analyzed the influence of the expert and non-expert annotators, and evaluated model fairness, specifically assessing differences in performance outcomes based on the speakers’ gender. We compared the results with the widely used IEMOCAP [14] and RAVDESS [15] datasets.
The main contributions of the article are: (1) presenting EMOVOME, the first public dataset with spontaneous real-life emotions for SER, and making it available to the scientific community; (2) analyzing the SER performance of the state-of-the-art pre-trained speech models in real-life spontaneous conversations, and comparing the results with two non-natural reference datasets; (3) analyzing the influence of EMOVOME annotators on SER results; and (4) evaluating models’ gender fairness. This study emphasizes the disparity between staged or laboratory settings and real-life scenarios, aiding in the advancement of recognizing authentic emotions.
The paper is organized as follows. Section 2 reviews existing SER datasets and related works. Section 3 describes EMOVOME and Section 4 other datasets used. Section 5 explains the methodology for creating SER models, and Section 6 reports the results. Section 7 discusses the results, and the paper concludes with a summary and future directions.
@@ II Related work
II-A Spontaneous datasets for SER
The availability of spontaneous datasets in the literature is limited. We conducted a search using KAPODI [16], a recently published searchable database of emotional stimulus sets, along with a review of the relevant literature. Table I presents the spontaneous datasets identified for English and Spanish, indicating the recording environment (RE) and its details, and whether they are free to access.
TABLE I: Spontaneous datasets for SER in English and Spanish. The asterisk (*) indicates if the dataset is multilingual.
Dataset RE Details Free
English
AFEW [17] Staged TV shows Yes
Belfast Naturalistic [18] Staged TV shows Yes
Castaway Reality TV [19] Staged TV shows Yes
MSP-PODCAST [20] Staged Podcasts Yes
AVIC [21] Lab Guided interaction Yes
NATURAL [9] Real-life Call-center (human) No
Lee & Narayanan [10] Real-life Call-center (machine) No
SUSAS [11] Real-life Pilot communications Yes
Doctor-patient interview No
Spanish
MOUD [12] Staged YouTube videos Yes
CMU-MOSEAS * [13] Staged YouTube videos Yes
For English, we identified four [17, 18, 19, 20] spontaneous datasets recorded in staged scenarios, primarily featuring clips from TV shows, and one [21] recorded in a laboratory setting, capturing interactions during a simulated product presentation. Regarding real-life scenarios, NATURAL [9] and Lee & Narayanan [10] datasets include recordings from call-centers, with interactions involving a human agent and a machine agent, respectively. Both datasets primarily contain neutral speech, resulting in highly unbalanced data. Additionally, unlike previous datasets, those are restricted to public use. SUSAS dataset[11] contains two real-life sets: isolated words from aircraft communications and doctor-patient interviews, but only the former is commercially available.
For Spanish, we found only two datasets in the literature: MOUD [12] and CMU-MOSEAS [13]. Both datasets are publicly available; however, they primarily consist of monologue clips sourced from YouTube videos, recorded in staged settings. Additionally, the original clips in CMU-MOSEAS are not disclosed, with only high-level features provided. The MOUD dataset presents another limitation, with the majority of speakers being female (84 out of 105), while the CMU-MOSEAS dataset does not specify gender distribution, making it challenging to develop gender-fair SER models.
II-B Performance of SER models
Research on SER using Spanish datasets is limited. For this reason, we reviewed prior studies that utilized Spanish datasets, including non-spontaneous ones, focusing on adult speech. Our analysis identified twelve Spanish datasets. Table II summarizes key details: expression authenticity (acted, elicited, or spontaneous), recording environment (RE), number of samples (N), number of speakers (Spk) with gender distribution (males/female, M/F), emotion model (with the number of categories for discrete models), and free use. For the four multilingual datasets (marked with an asterisk), only the Spanish partition is detailed. Additionally, we report the classification performance of SER models from the literature for each Spanish datasets, using two evaluation metrics: Weighted Accuracy (WA) and Unweighted Accuracy (UA). WA, or simply “accuracy”, measures the overall proportion of correct predictions but can be biased toward majority classes in imbalanced datasets. UA, commonly used in SER studies, provides a fairer evaluation by averaging accuracy across all classes. The results indicate if a speaker-dependent (SD) or speaker-independent (SI) approach was used, i.e. whether same or different speakers were used for training and testing (if reported in the source articles). Multilingual and cross-lingual models are excluded from the table, as they pose greater challenges, and monolingual models typically achieve better performance when trained on comparable sample sizes [22].
TABLE II: Datsets for SER in Spanish and corresponding results for SER models. An asterisk (*) denotes multilingual datasets.
Dataset RE N Spk (M/F) Emotion Free Models results
UA (%) WA (%)
Acted
Iriondo et al. [23] Lab 336 8 (4/4) Discrete (7) No - -
Spanish Emotional Speech [24] Lab 1288 1 (1/0) Discrete (4) No - -
Martínez & Cruz [25] Lab 380 $>$15 (-/-) Discrete (5) No - 55 [25]
Emotional Mexican Spanish speech [26] Lab 240 6 (3/3) Discrete (4) No $\geq 95$ [26] -
Spanish Expressive Voices [27] Lab 3890 2 (1/1) Discrete (7) No - 95 [27]
RekEmozio * [28] Lab 2618 10 (5/5) Discrete (7) No - SI: 74.82 [29]
Mexican Emotional Speech Database [30] Lab 3456 8 (4/4) Discrete (6) Yes 93.90 / 89.49 M/F [30] -
Emotional speech synthesis [31]
(from INTERFACE * [32])
Lab 5520 2 (1/1) Discrete (7) Yes 89.4 / 90.9 M/F [33]
91.16 [34], 91.51
[35], 94.01 [36]
EmoFilm * [37] Staged 342 57 (33/24) Discrete (5) Yes SI: 69.15 [22]
SI: 70.76 [22]
@@ II-B Performance of SER models
Spontaneous
MOUD [12] Staged 550 105 (21/84)
Discrete (3)
from valence
Yes -
46.75 (positive
@@
CMU-MOSEAS * [13] Staged 10000 341 (-/-) Discrete (6) Yes - -
SER performance is conditioned by the type of data used. For acted datasets, the table shows WA and UA values equal to or exceeding 85% (e.g. [26, 27, 30, 33, 34, 35, 36, 38]). However, performance drops to 70 to 75% [22, 29] when SI models are used, an approach which is known to yield lower results compared to SD models [5]. An exception is seen in [25], where an accuracy of 55% was achieved for the classification of five emotions. The model was trained on actors simulating emotions while reading sentences and tested on movie clips, leading to a domain mismatch that likely caused the significant performance drop. Lower results on SI models also occur in the elicited EmoSpanishDB dataset [39], where seven emotions where classified with UA scores of 56.3% and 42.1% for SD and SI, respectively. A refined version of the dataset, EmoMatchSpanishDB, improved performance by removing mislabeled samples that did not align with the intended elicited emotions, achieving UA scores of 65.0% (SD) and 64.2% (SI). Finally, the spontaneous dataset shows the lowest performance. In [12], a classification accuracy of 46.75% was achieved for differentiating between positive and negative samples, highlighting the challenges of spontaneous emotional data in SER.
@@
Regarding the techniques used to create the SER models, a noteworthy paradigm shift involves leveraging large pre-trained models in emotion recognition tasks. The study in [22] used the pre-trained wav2vec2-large-robust model fine-tuned on MSP-Podcasts dataset from [40] to extract speech embeddings. These embeddings were then fed into a support vector machine, obtaining 69.15% unweighted accuracy for five emotion categories on the EmoFilm dataset. To our knowledge, this is the only previous study using pre-trained models for monolingual SER models in Spanish. Nevertheless, an increasing number of recent publications have focused on the use of pre-trained models with English datasets for emotion categories and less frequently for dimensions.
@@
As for studies focused on discrete emotion models, a very recent study [40] summarized the state-of-the-art results for 4-class emotion classification in the widely used IEMOCAP dataset, including studies using the pre-trained wav2vec 2.0 and HuBERT models. Comparing the cross-validation results, the UA values range from 60.0% to 74.3%, while the weighted average recall values vary from 62.6% to 79.6%, both top models. Overall, HuBERT surpassed the performance of the wav2vec 2.0 models. Other works have explored several pre-trained models and emotional datasets. In [41], they used eight pre-trained models (including wav2vec 2.0) and four datasets to predict 6-7 emotions. The results showed that pre-trained models for speaker recognition (x-vector and ECAPA) achieved the highest scores, possibly due to their learned ability to identify unique features within an individual’s speech. Subsequently, UniSpeech-SAT achieved the best results, a model pre-trained using multitask learning, including the speaker identity. In another study [42], the authors compared nineteen pre-trained models (including wav2vec 2.0, HuBERT, UniSpeech-SAT and wavLM) and five datasets to predict 4-6 emotion categories using a SI approach. They found that the best scores were achieved with WavLM, UniSpeech-SAT, and HuBERT, all three in the large version.
@@
As for studies on continuous emotion models, [43] used a multimodal model (audio + text) as a teacher to fine-tune HuBERT embeddings to predict valence, arousal and dominance on MSP-Podcast dataset [44]. They obtained state-of-the-art Concordance Correlation Coefficient (CCC) values of 0.757, 0.627 and 0.671 for arousal, valence and dominance, respectively. The previous state-of-the-art performance for valence was 0.377 [45], so 0.627 represents a substantial improvement. They also replicated the results for the IEMOCAP dataset, achieving again state-of-the-art CCC results for valence (0.667), arousal (0.582) and dominance (0.545). In [40], the authors conducted an exhaustive analysis using different variants of the pre-trained wav2vec 2.0 and HuBERT models for valence, arousal and dominance prediction on MSP-Podcast. They obtained the best result in the literature for valence prediction using only audio, with a wav2vec2-large-robust model that achieved a CCC of 0.638. Notably, their results showed that data used for pre-training the models and the fine-tuning of the transformer layers had a strong influence on valence prediction. Both factors played a role in shaping the models’ ability to implicitly incorporate linguistic information embedded in the audio signal. This, in turn, accounts for their success in valence prediction, achieving similar performance to multimodal models that integrate explicit textual information.
II-C Fairness evaluation